# Exploring Power Side-Channel Challenges in Embedded Systems Security

### v1

@@ 4.2 PSC Applications
  • Control Flow Graph: As mentioned in the previous subsection, the PSC leaks information about the instructions being executed in the program. As a result, one can use this information to construct the Control Flow Graph (CFG) of the application and monitor its different run-time behaviors. For example, CFG data can be used to recognize abnormal patterns to detect intrusion (i.e., Intrusion Detection System) [12, 11], or it can be used to check the program’s integrity [9]. Here, the implemented model should be lightweight enough to operate in real-time, especially in critical industrial applications [32].
@@
  • Fuzzing: Fuzzing has been recognized as one of the most efficient and practical methods to find security vulnerabilities in both hardware and software. Considering that the PSC leaks information about the CFG, it can be used to guide a fuzzer. Recently, PSC has been used as a coverage feedback mechanism to guide a fuzzer in a theoretical way [10]. Detecting conditional branch instructions is crucial in coverage-guided fuzzing because they often signal the possibility of new basic blocks and, consequently, new execution paths. This information collected from PSC can guide a fuzzer in generating inputs that explore a broader portion of the program, helping to identify possible security vulnerabilities across the program’s entire space. Meanwhile, crash detection and root cause analysis in embedded devices is one of the open problems in embedded fuzzing. The existing challenges regarding embedded fuzzing have already been discussed in [33].
@@
  • Software Composition Analysis: PSC can enable software composition analysis on embedded devices without requiring firmware access. By analyzing power consumption patterns during execution, it may be possible to identify the presence and behavior of known third-party software components. Since different software modules exhibit distinct power usage profiles, these patterns can help determine which libraries or components are running.
@@
  • Peripheral Interaction: Another application of the PSC in embedded devices is detecting peripheral interactions. Generally, a microprocessor communicates with external devices using its various communication interfaces, such as UART, SPI, and Ethernet. These devices can be either on the PCB, such as an SPI flash, or they can be external devices with a separate PCB, such as a slave processor or sensor. In general, the power consumption for these communications is significantly more than normal internal processes as these interfaces communicate with external devices. Therefore, if the attacker can monitor the power consumption, then it is possible to decode the data that is being transferred using these interfaces. In the case that the external device is on the same PCB (the SPI flash example), the pins might not be accessible for decoding the data lines because of the BGA packaging. While the attacker can directly monitor the power pins from different locations and decode the data.
@@ 4.3 Experimental Setup
For each of the challenges discussed in the following sections, we designed experiments to demonstrate them in the context of embedded system security. We used a Tektronix MSO66B oscilloscope equipped with passive probes that offer a 1 GHz bandwidth. The sampling frequency varied depending on the experiment, ranging from 50 MS/s to 3.125 GS/s. Additionally, this oscilloscope provides the capability to design custom filters, which we utilized by applying a high-pass filter as detailed in Section 6.
For our experiments, we selected five different targets to cover different scenarios in PSC measurements:
@@
  • General-Purpose Development Board (Nucleo)): A Nucleo-144 [34] board, which is a general-purpose development board, and its design does not consider the PSC attacks, such as a shunt resistor. This board features a Cortex-M4 STM32F429 microcontroller, an onboard ST-LINK programmer and debugger, and an Ethernet socket for network communications.
  • Real-World Black-Box Target (UBlox): A UBlox ZED-F9P-02B GNSS receiver [35] is used as a real-world black-box embedded target. This device’s firmware and Instruction Set Architecture (ISA) are not available, and only the interface semantics (protocol specifications) are available, making it a suitable target for testing PSC analysis in scenarios with limited information.
  • Real-World Black-Box Target (Satellite): Satellite On-Board RF-Transceiver. We employ this device to evaluate the use of PSC for observing inaccessible data busses (e.g., busses on internal layers of a multi-layer PCB) non-destructively. Other than what can be identified from the markings of chips on the PCB, the circuits and the device’s firmware are entirely unknown.
@@ 4.4 Probing
In the embedded system security domain, selecting a probing point is significantly more challenging than in other domains. For example, if a researcher wants to test cryptographic algorithms against PSC attacks, these algorithms will be implemented on ideal hardware specifically designed for PSC measurements (such as [36, 37]) with an ideal probing point included. Similarly, in defensive scenarios such as IDS, the defender could enable the inclusion of optimal probing points for more reliable PSC analysis in the PCB design stage. However, if the defender wants to deploy the PSC-based IDS to an existing embedded device, getting an idea probing point would not be easy. Moreover, obtaining such an ideal probing point is even harder in the embedded system security attack scenario. Typically, attackers are dealing with an existing embedded device where they aim to deploy an attack. In these cases, provisions such as a shunt resistor for PSC measurements are absent, forcing the attacker to either add these requirements (often impractical) or resort to a noisy probing point on the device’s main power supply.
@@
To show the impact of noise on the probing point, we design an experiment. We choose two targets: the first one is the ideal target, and the second one is the STM32F4 development board with a shunt resistor manually attached to the main power supply lines of the board. We ran a program with random combinations of four instructions—ADD, SUB, AND, and ORR—on both boards and recorded the power traces. The sampling frequency was set to 1.25 GS/s, and the bandwidth was 20 MHz, satisfying the Nyquist criterion. For classification, we used a Multi-Layer Perceptron (MLP), a method previously employed for PSC disassembly in [29]. The confusion matrix is shown in Figure 2. On the target with the soldered shunt resistor, the classification accuracy was 93%, while on the development board, the accuracy dropped to 51%. This result demonstrates that the presence of other electrical components in the power introduces significant noise, which drastically reduces the accuracy of instruction profiling when using the same method.
@@ 4.6 Signal Processing
As previously discussed, various signal processing techniques are commonly used in traditional PSC attacks, such as those targeting cryptographic algorithms, to enhance the Signal-to-Noise Ratio (SNR) and improve attack efficiency. However, in the context of embedded firmware, which often includes an embedded RTOS, additional background processes introduce extra noise that complicates the signal processing. This noise is primarily timing-related, causing misalignment in the power traces. It is important to note that this misalignment is unintentional and differs from the intentional misalignment introduced by operation shuffling in cryptographic algorithms. In applications like fuzzing and Intrusion Detection Systems (IDS), the most prevalent method is based on pattern matching [32, 10]. However, variations in background processes from one execution to another can lead to misalignment issues. This challenge is thoroughly demonstrated in [32], where the authors aim to detect intrusions. They propose a solution called dynamic window matching to address this problem, but their approach still relies on pattern matching, which involves a heavy computational effort to search through the entire power trace for specific patterns. Moreover, if profiling the target is not feasible, finding the correct point to align the traces becomes impractical.
@@
With the advancement of Machine Learning (ML) and Deep Neural Networks (DNNs), these techniques have proven to be highly effective tools for signal processing. They have been applied to various PSC applications, serving either as classifiers or as pure signal processing tools for different purposes. For example, Convolutional Neural Networks (CNN) have been used as a classifier for breaking cryptographic algorithms [38, 39]. Additionally, they have been used for power trace alignment [40, 41]. Furthermore, autoencoders as an unsupervised ML model are used for noise reduction of power traces [42, 43]. Although ML-based models are efficient, they have a significant weakness in most embedded system security applications. ML methods are data-driven, meaning their performance depends heavily on the quality and representativeness of their training datasets. However, in applications where profiling is not feasible or where a dataset representing semi-real-world power traces is unavailable, these models struggle to generalize effectively, and the efficiency of ML-based methods can decrease drastically, hence limiting their applicability in such scenarios.
@@
Moreover, ML models typically assume that the training and evaluation datasets follow an Identical and Independent Distribution (IID). This means that, for the ML model to perform effectively during an attack, the attacker must have access to the same firmware (or at least firmware with a similar distribution of instructions) during the training phase. If this condition is not met, the trained ML model may not perform well in the attack phase. The issue of differing distributions between the training and evaluation datasets is known as Covariate Shift in the ML domain, and it remains an unresolved problem. All the studies mentioned in Table II use datasets with the same distribution during both the profiling and attack phases, thereby avoiding the covariate shift problem. However, this also means that these attacks are not applicable in scenarios where the attacker does not have access to the firmware. As a result, none of the existing PSC-based disassemblers can effectively operate in black-box scenarios, rendering them inefficient in real-world cases where such access is unavailable.
@@
To illustrate the challenge of covariate shift, we developed a mini prototype random assembly code generator, similar to the approach proposed in [30], but this time tailored for the ARM Cortex-M4. Using this tool, we generated five random programs with uniformly distributed instructions. Additionally, we selected the AES128 encryption algorithm from tiny-AES-C [44] as a real-world program. We conducted our experiments on the ideal STM32F3 target, executing all six programs (the five random ones and the AES128) while collecting power traces.In a manner similar to the experiment described in Section 4.4, we trained an MLP classifier with two hidden layers on four of the five random programs. We then evaluated the classifier twice: once on the fifth random program and once on the AES128 program. The accuracy dropped significantly from 81% to 41%, demonstrating the impact of covariate shift. The drastic drop in accuracy during the second experiment occurred because the distribution of instructions in the AES encryption algorithm is not uniform, highlighting the covariate shift problem. In the Appendix, Figure 13 displays the confusion matrix for these two scenarios.
@@ 4.7 Run-time Overhead
As discussed earlier, repeating measurements for a single input is a common method to reduce noise in PSC. However, in real-time applications such as fuzzing or intrusion detection, this practicability of the approach is limited due to the time-consuming nature of multiple executions and the requirement to reset the target to the same state each time. Consequently, the number of measurements that can be taken in real-time applications is a critical factor. In the domain of PSC-guided fuzzing, we only find one notable study, [10], which implements fuzzing in a theoretical manner using a white-box scenario. This study is considered theoretical because it relies on previously recorded power traces instead of recording power traces and guiding the fuzzer in real-time. Additionally, even if power profiling is feasible in a white-box scenario, it is not possible in black-box systems. Moreover, the paper does not provide details on execution time or a comparison with a blind fuzzer. This omission is significant because if a PSC-assisted fuzzer is slower than a blind fuzzer, its use may not be time-efficient. As a result, the challenge of effectively using PSCs in real-time applications such as fuzzing remains unaddressed, highlighting a gap in current research.
@@ 4.8 Application Specific Challenge
Apart from the mentioned challenges, there are a few application-specific challenges in the PSC. Specifically, we would like to discuss crash detection and crash clustering challenges in fuzzing application.
Crash Detection and Crash Clustering in Fuzzing
In fuzzing, crash detection is crucial. Firmware reboot is a common crash signal but crashes may not result in a reboot instantly. Instead, a fault handler is typically invoked to manage the unexpected behavior. Faults can vary, including CPU faults and software faults, consequently, different embedded devices may run firmware on either bare metal or an RTOS, leading to varied implementations of fault handlers. This variability makes profiling even more challenging. Additionally, identifying the specific fault handler involved can provide valuable insights into the type of crash and its root cause. However, crash detection in embedded devices differs significantly from other fuzzing applications. In these scenarios, particularly in black-box environments, there is often no debugging interface available, making crash detection and root cause analysis a rigorous task.
@@ 5.1 Control Flow Fingerprinting Using PSC
PSC was previously used to gather information about the code execution [32], but the mentioned challenges affect the measurements and reduce the accuracy in different applications; most of them are unsuitable for non-profiling scenarios. In this section, we propose our algorithm based on wavelet transform to distinguish the different code executions from power traces. As discussed in Section 4.6, using a ML model for signal processing is not desired here as they suffer from the covariate shift problem and are not suitable for black-box applications.
Before explaining the details of the algorithm, a discussion about what is actually going to be fingerprinted is required. The typical power consumption of a code execution depends on the instructions and the data. Among instructions, branch instructions are the more interesting ones as if the branch is taken; it means that not only the current code execution should stop, but also the pipeline should be cleared, and the program counter (PC) should change. These actions leave a distinguishable pattern in power trace as the amount of consumed power changes drastically. Consequently, these drastic changes lead to high-frequency changes in power traces. Therefore, power consumption leaks information about the CFG and can be used to track the code execution. However, PSC signals are noise-sensitive, and typically, multiple measurements (can vary from ten to thousands) [45] are required to reduce the noise. This raises challenges in using PSCs in real-time applications.
@@ 5.1.1 Frequency Domain Fingerprinting
Target Accuracy False positive Branches
AES 128 71.29% 9.25% 586/862
SHA 256 75.27% 14.86% 633/841
@@
With the identified source of information leakage for CFG fingerprinting, we now introduce our signal-processing algorithm. Assuming that $T(t)$ is a measured raw power trace, first, a wavelet transform is applied to the power trace to extract the frequency domain representation. Then, the locations corresponding to high-frequency changes should have higher values and can be filtered by choosing a threshold $\Theta$:
$\Omega^{\prime\prime}\ =\ \operatorname*{argmax}_{t}\ |W(T(t))|\ >\ \Theta$
Where $\mathrm{W}$ is the wavelet transform and $\mathrm{\Omega^{\prime\prime}}$ is the possible location for unusual changes in the PC. Also, $\Theta$ is defined as follows:
$\Theta\ =\ \theta\ *\ \max{|W(T(t))|}$
Where $\theta$ is a threshold parameter that is set by the user. Here, $\mathrm{\Omega^{\prime\prime}}$ may have duplicate values for the same location because these high-frequency changes are observable in multiple frequencies as the wavelet transform filters the signal in a range of frequencies. We call these duplicate values spectral duplicates. Therefore, we need to eliminate these duplicate values.
@@
Figure 4: Comparison of the extracted fingerprint with our proposed method for two different algorithms.
where $\mathrm{\Omega^{\prime}}$ is a set of candidates representing locations for severe changes in PC. To turn $\mathrm{\Omega^{\prime}}$ to a vector that fingerprints the program flow, we need to transform these values to instruction locations. Hence, two different approaches can be used: 1) clock reconstruction or 2) division by sample per clock. Clock reconstruction provides a higher level of accuracy. However, assuming the attacker has access to the clock signal is a strong assumption. If the processor uses PLL to modify the external crystal’s clock frequency, calculating the PLL multiplier values becomes impossible in a black-box system. In this case, we take the second approach and assume that the attacker can estimate the number of samples per clock, which is a reasonable assumption when using a power trace.
As a result, the values in $\mathrm{\Omega^{\prime}}$ are divided by the number of samples per clock. Here, another step is applied to eliminate the duplicate values that may exist due to having different samples from the same clock cycle in $\mathrm{\Omega^{\prime}}$. We call these duplicate values temporal duplicates. The final set of values that are represented as a vector $\mathrm{\Omega}$ is a unique identifier of the possible locations for severe changes in the PC.
@@
The most important advantage and a major contribution of our algorithm is that the proposed method does not need any information about the target firmware implementation, as we do not need to perform profiling. To demonstrate the practicality of our algorithm, we will show that an attacker can use it to extract information from a single execution run (without profiling), making it a proper choice for real-time applications.
@@
To evaluate the efficiency of our wavelet-based control flow fingerprinting algorithm with a single execution, we must first test our concept on a real-world microcontroller. Also, a thorough assessment of the accuracy and false positives is required. Therefore, we choose two different programs, namely tinyaes128 [44] and tinysha256 [46], and assess our algorithm’s efficiency on them. For all three target programs, the threshold $\mathrm{\theta}$ is set to 0.3, the sampling frequency was 3.125 GS/s, and the target processor runs on an 8 MHz clock. We ran these programs on an STM32F3 board with a soldered shunt resistor.
The result of our experiment is illustrated in Table III. Accuracy is the number of branches detected compared to the existing branches, and false positive is the rate of falsely detected branches compared to the existing branches. As the results indicate, the proposed method can identify changes in the PC and represent it as a unique fingerprint for the code execution. Figure 4 shows the differences between the extracted vectors for the discussed programs. As can be seen, the ten rounds of AES128 are observable. Also, the differences between the execution paths of the programs show our method’s different output for different behaviors on the PC.
We argue that our proposed signal processing algorithm performs well in fingerprinting the control flow of a black-box target without any profiling. The most crucial point about our algorithm is that it can work with a single power trace. However, the downside of this method is the high resolution (3.125 GS/s) and high amount of computations. Also, as mentioned in Section 4.5, peripheral access can cause massive voltage peaks that may affect the accuracy of our method. Later, we will evaluate our method in such cases and try to improve it.

### v2

@@ 4.2.2 Control Flow Graph:
As mentioned in the previous subsection, the PSC leaks information about the instructions being executed in the program. As a result, one can use this information to construct the Control Flow Graph (CFG) of the application and monitor its different run-time behaviors. For example, CFG data can be used to recognize abnormal patterns to detect intrusion (i.e., IDS) [32, 11], or it can be used to check the program’s integrity [9]. Here, the implemented model should be lightweight enough to operate in real-time, especially in critical industrial applications [12]. The attacker model of the mentioned intrusion detection systems are white-box as they have access to the ISA and firmware of the target.
4.2.3 Fuzzing:
@@
Fuzzing has been recognized as one of the most efficient and practical methods to find security vulnerabilities in both hardware and software. Considering that the PSC leaks information about the CFG, it can be used to guide a fuzzer. Recently, PSC has been used as a coverage feedback mechanism to guide a fuzzer in a theoretical way [10]. The attacker model in this work is a white-box model as they have access to the ISA and firmware in profiling phase, and they use the same setting in the attack phase. However, this attacker model does not generalize to real-world scenarios as usually the attacker does not have ISA and/or firmware. Detecting conditional branch instructions is crucial in coverage-guided fuzzing because they often signal the possibility of new basic blocks and new execution paths. This information collected from PSC can guide a fuzzer in generating inputs that explore a broader portion of the program, helping to identify possible security vulnerabilities across the program’s entire space. Meanwhile, crash detection and root cause analysis in embedded devices is one of the open problems in embedded fuzzing. The existing challenges regarding embedded fuzzing have already been discussed in [17], therefore we skip it here.
4.2.4 Software Composition Analysis:
@@
PSC can enable software composition analysis on embedded devices without requiring firmware access. By analyzing power consumption patterns during execution, it may be possible to identify the presence and behavior of known third-party software components. Since different software modules exhibit distinct power usage profiles, these patterns can help determine which libraries or components are running.
4.2.5 Peripheral Interaction:
@@
Another application of the PSC in embedded devices is detecting peripheral interactions. Generally, a microprocessor communicates with external devices using its various communication interfaces, such as UART, SPI, and Ethernet. These devices can be either on the PCB, such as an SPI flash, or they can be external devices with a separate PCB, such as a slave processor or sensor. In general, the power consumption for these communications is significantly more than normal internal processes as these interfaces communicate with external devices. Therefore, if the attacker monitors the power consumption, then they can decode the data that is being transferred using these interfaces. In the case that the external device is on the same PCB (the SPI flash example), the pins might not be accessible for decoding the data lines due to the BGA packaging. While the attacker can directly monitor the power pins from different locations and decode the data.
@@ 4.3 Experimental Setup
We design an experiment for each challenge we introduce to demonstrate it in the context of embedded system security. We use a Tektronix MSO66B oscilloscope equipped with passive probes that offer a 1 GHz bandwidth (Appendix 8.1). The sampling frequency varies depending on the experiment and satisfying Nyquist criterion, ranging from 25 MS/s to 3.125 GS/s.
For our experiments, we select five different targets to cover different scenarios in PSC measurements:
@@
  • General-Purpose Development Board (Nucleo): A general-purpose development board, Nucleo-144 [33], without any PSC measurement considerations, such as a shunt resistor. This board features a Cortex-M4 STM32F429 microcontroller, an onboard ST-LINK programmer and debugger, and an Ethernet socket for network communications.
  • Real-World Black-Box Target (UBlox): A UBlox ZED-F9P-02B GNSS receiver [34] is used as a real-world black-box embedded target. This device’s firmware and ISA are unavailable, and only the interface semantics (protocol specifications) are available, making it a suitable target for testing PSC analysis in scenarios with limited information.
  • Real-World Black-Box Target (Satellite): We employ a satellite on-board RF-transceiver to evaluate the application of PSC for observing inaccessible data busses (e.g., busses on internal layers of a multi-layer PCB) non-destructively. Other than what can be identified from the markings of chips on the PCB, the circuits and the device’s firmware are entirely unknown.
  • Real-World High-end Target (Raspberry Pi): We use a Raspberry Pi Zero 2 W [35] board as a high-end target (with a 64-bit ARM Cortex-A53 processor clocked at 1GHz) running a raspbian kernel. This board has a processor power test pin, which we use for power measurement.
@@ 4.4 Probing
In the embedded system security domain, selecting a probing point is significantly more challenging than in other domains. For example, in cryptographic attacks, the attacker uses an ideal board designed for PSC measurements (such as [13, 36]). Similarly, in defensive scenarios such as IDS, the defender could enable the inclusion of optimal probing points for more reliable PSC analysis in the PCB design stage. However, if the defender wants to deploy the PSC-based IDS to an existing embedded device, finding an ideal probing point would requires hardware manipulation. Typically, attackers are dealing with an existing embedded device where they aim to deploy an attack. In these cases, provisions such as a shunt resistor for PSC measurements are absent, forcing the attacker to either add these requirements (often impractical) or resort to a noisy probing point on the device’s main power supply.
@@
To show the impact of noise on the probing point, we design an experiment. We choose two targets: the first one is the ideal target, and the second one is the STM32F4 development board with a shunt resistor manually attached to the main power supply lines of the board. We run a program with random combinations of four instructions—ADD, SUB, AND, and ORR—on both boards and record the power traces. The sampling frequency is set to 1.25 GS/s, and the bandwidth is 20 MHz. For classification, we use a Multi-Layer Perceptron (MLP), a method previously employed for PSC disassembly in [29]. On the target with the soldered shunt resistor, the classification accuracy is 93%, while on the development board, the accuracy drops to 51% (Figure 14 in Appendix 8.2). This result demonstrates that the presence of other electrical components on the PCB introduces significant noise, which drastically reduces the accuracy of instruction disassembly when using the same method.
@@ 4.6 Signal Processing
Traditional PSC attacks use various signal processing techniques to enhance the Signal-to-Noise Ration (SNR) and improve attack efficiency. However, in the context of embedded firmware, which often includes an embedded RTOS, additional background processes introduce extra noise that complicates the signal processing. This noise is primarily timing-related, causing misalignment in the power traces. It is important to note that this misalignment is unintentional and differs from the intentional misalignment introduced by operation shuffling in cryptographic counter measures. In applications such as fuzzing and IDS, the most prevalent method is based on pattern matching [12, 10]. However, variations in background processes from one execution to another can lead to misalignment issues. This challenge is thoroughly demonstrated in [12], where the authors aim to detect intrusions. They propose a solution called dynamic window matching to address this problem, but their approach still relies on pattern matching, which involves a heavy computational effort to search through the entire power trace for specific patterns. Moreover, if profiling the same target is not feasible, finding the correct point to align the traces becomes impractical.
@@
With the advancement of ML and Deep Neural Networks (DNNs), these techniques have proven to be highly effective tools for signal processing. Researchers apply them to various PSC applications, either as classifiers or as pure signal processing tools for different purposes. For example, researchers use Convolutional Neural Networks (CNN) as a classifier for breaking cryptographic algorithms [37, 38]. Additionally, they have been used for power trace alignment [39, 40]. Furthermore, autoencoders as an unsupervised ML model are used for noise reduction of power traces [41, 42]. Although ML-based models are efficient, they are data-driven. Meaning their performance depends heavily on the quality and representativeness of their training datasets. However, in applications where profiling is not feasible or where a dataset representing semi-real-world power traces is unavailable, these models struggle to generalize effectively limiting their applicability in such scenarios.
@@
Moreover, ML models assume that the training and evaluation datasets follow an Identical and Independent Distribution (IID). This means that, for the ML model to perform effectively during an attack, the attacker must access to the same firmware (or firmware with a similar distribution of instructions) during the training phase. If this condition is not met, the trained ML model may not perform well in the attack phase. The issue of differing distributions between the training and evaluation datasets is known as Covariate Shift in the ML domain, and it remains an unresolved problem. All the studies mentioned in Table II use datasets with the same distribution during both the profiling and attack phases, thereby avoiding the covariate shift problem. However, these attacks are not applicable in scenarios where the attacker cannot access to the firmware. As a result, none of the existing PSC-based disassemblers can effectively operate in black-box scenarios, rendering them inefficient in real-world cases where such access is unavailable.
@@
To illustrate the challenge of covariate shift, we developed a simple random assembly code generator, similar to the approach proposed in [30], but this time tailored for the ARM Cortex-M4. Using this tool, we generate five random programs with uniformly distributed instructions. Additionally, we select the AES128 encryption algorithm from tiny-AES-C [43] as a real-world program. We conduct our experiments on the ideal STM32F3 target, executing all six programs (the five random ones and the AES128) while collecting power traces. In a manner similar to the experiment described in Section 4.4, we train a MLP classifier with two hidden layers on four of the five random programs. We then evaluate the classifier twice: once on the fifth random program and once on the AES128 program. The accuracy drops significantly from 81% to 41%, demonstrating the impact of covariate shift. The drastic drop in accuracy during the second experiment occurres because the distribution of instructions in the AES encryption algorithm is not uniform, highlighting the covariate shift problem. In the Appendix 8.3, Figure 15 displays the confusion matrix for these two scenarios.
@@ 4.7 Run-time Overhead
Repeating measurements for a single input is a common method to reduce noise in PSC attacks. However, in real-time-constrained applications such as fuzzing or intrusion detection, the practicability of the approach is limited due to the time-consuming nature of multiple executions and the requirement to reset the target to the same state each time. Therefore, execution per second is a limiting factor. In the domain of PSC-guided fuzzing, we only find one notable study, [10], which theoretically implements fuzzing in a white-box scenario. We consider this study theoretical because it relies on previously recorded power traces instead of recording power traces and guiding the fuzzer in real-time. Additionally, even if power profiling is feasible in a white-box scenario, it is not possible in black-box systems as we do not have the firmware as the ground truth. Moreover, the paper does not provide details on execution time or a comparison with a blind fuzzer. This omission is significant because if a PSC-assisted fuzzer is slower than a blind fuzzer, its use may not be time-efficient. As a result, the challenge of effectively using PSCs in real-time applications such as fuzzing remains unaddressed, highlighting a gap in current research.
@@ 4.8 Application Specific Challenge
Apart from the mentioned challenges, there are a few application-specific challenges in the PSC. Specifically, we discuss crash detection, crash clustering, and cross-architecture generalization in fuzzing applications.
4.8.1 Crash Detection and Crash Clustering in Fuzzing:
In fuzzing, crash detection is crucial. A reboot is a common crash signal this typically involves a fault handler that manages unexpected behavior. Faults vary, including CPU and software faults. Different embedded devices may run bare-metal-firmware or an RTOS, leading to varied implementations of fault handlers. This variability makes profiling even more challenging. Identifying the specific fault handler involved can provide valuable insights into the type of crash and its root cause. In these scenarios, particularly in black-box environments, there is often no debugging interface available, making crash detection and root cause analysis a rigorous task.
4.8.2 Cross Architecture CFG Reconstruction:
Another significant challenge for CFG analysis or IDS applications is their effectiveness across different architectures. Profiling-based approaches face a major hurdle in generalizing across architectures, as various architectures exhibit distinct leakage patterns due to differences in design, such as the number of pipeline stages. Most existing research has evaluated their methods on a single target or within a specific family of targets (as seen in Table II). Consequently, the challenge of evaluating these methods across different architectures remains an open issue for CFG and IDS-related applications.
Figure 3: Software Noise vs Hardware Noise
4.9 Hardware noise vs. Software noise
In this section, we discussed the challenges of using PSC in system security. From a high-level overview, we group the source of these challenges in two subgroups (Figure 3).
  • Hardware Noise: This includes the noise generated by the PCB, peripherals, and components on it.
  • Software Noise: This includes system-level interferences such as other processes, interrupts, and system calls. Software noise is the major challenge in using PSC in system security applications.
@@ 5.1 Control Flow Fingerprinting Using PSC
PSC was previously used for code execution tracking [12], but the mentioned challenges affect the measurements and reduce the accuracy in different applications; most are unsuitable for cross-device scenarios. In this section, we propose our algorithm based on wavelet transform to distinguish different code executions from power traces. As discussed in Section 4.6, using ML models for signal processing is not desired here as they suffer from the covariate shift problem and are unsuitable for black-box applications.
Before explaining the details of the algorithm, we discuss what we will fingerprint. The typical power consumption of a code execution depends on the instructions and the data. Among instructions, branch instructions are the more interesting ones as if the branch is taken; not only should the current code execution stop, but the pipeline should also be cleared, and the program counter (PC) should change. These actions leave a distinguishable pattern in power trace as the amount of consumed power changes drastically. Consequently, these drastic changes lead to high-frequency changes in power traces. Therefore, power consumption leaks information about the CFG and can be used to track the code execution. However, PSC signals are noise-sensitive, and typically, multiple measurements (varies from ten to thousands) [44] are required to reduce the noise. This raises challenges in using PSCs in real-time applications.
@@ 5.1.1 Frequency Domain Fingerprinting
Target Accuracy False positive Drastic PC changes
AES 128 100% 10.82% 862
SHA 256 100% 11.30% 841
LZ4 100% 6.27% 2266
@@
With the identified source of information leakage for CFG fingerprinting, we now introduce our signal-processing algorithm. Assuming that $x(t)$ is a measured raw power trace, first, a wavelet transform is applied to the power trace to extract the frequency domain representation, and the result is a $n*m$ matrix:
$W_{x}(t,a)=\begin{bmatrix}g_{00}&g_{01}&.&g_{0m}\\ g_{10}&g_{11}&.&g_{1m}\\ .&.&...&.\\ g_{n0}&g_{n1}&.&g_{nm}\end{bmatrix}$
Where $a$ is the frequency scale of the wavelet. In this matrix, each row represents the wavelet transform for a fixed $a$ (daughter wavelet), and each column shows different scales in a single time slot. In this matrix, the values greater than a threshold $\Theta$ show the high-frequency changes with corresponding time slots. We filter the remaining values as follows:
$\Omega^{\prime\prime}\ =\ \operatorname*{argmax}_{t}\ |W_{x}(t,a)|\ >\ \Theta$
Where $\Omega^{\prime\prime}$ is a list of possible locations for high-frequency changes. Also, $\Theta$ is defined as follows:
$\Theta\ =\ \theta\ *\ \max{|W_{x}(t,a)|}$
Where $\theta$ is a threshold parameter that is set by the user. $\Omega^{\prime\prime}$ may have duplicate values for the same time slot because these high-frequency changes are observable in multiple frequencies as the wavelet transform filters the signal in a range of frequencies. We call these duplicate values spectral duplicates and they are usually repeated across the same time slot but in different rows (usually neighboring frequencies). Therefore, we eliminate these duplicate values:
@@
Figure 4: The fingerprint for AES128 in the presence of peripheral peaks, with ten rounds highlighted.
Where $\mathrm{\Omega^{\prime}}$ is a set of candidates representing time location (time slot) for high-frequency changes. To turn $\mathrm{\Omega^{\prime}}$ to a vector that fingerprints the program flow, we need to transform these values to instruction locations. Hence, we can use two approaches: 1) clock reconstruction or 2) division by sample per clock. Clock reconstruction provides a higher level of accuracy. However, assuming the attacker has access to the clock signal is a strong assumption. If the processor uses PLL to modify the external crystal’s clock frequency, calculating the PLL multiplier values becomes impossible in a black-box system. In this case, we take the second approach and assume that the attacker can estimate the number of samples per clock, which is a reasonable assumption when using a power trace. As a result, the values in $\mathrm{\Omega^{\prime}}$ are divided by the number of samples per clock. Here, another step is applied to eliminate the duplicate values that may exist due to having different samples from the same clock cycle in $\mathrm{\Omega^{\prime}}$. This mostly happens in cases where the sampling frequency is high enough that the size of the daughter wavelet $\psi_{a}(t)$ is similar to or smaller than a clock cycle. We call these duplicate values temporal duplicates. The final set of values that are represented as a vector $\mathrm{\Omega}$ is a unique identifier of the high-frequency changes in the signal $x(t)$.
@@
The most important advantage and a major contribution of our algorithm is that the proposed method does not need any information about the target firmware implementation, as we do not need to perform profiling on the same device. To demonstrate the practicality of our algorithm, we will show that an attacker can use it to extract information from a single execution run (without profiling the same device), making it a proper choice for real-time applications.
@@
To evaluate the efficiency of our wavelet-based control flow fingerprinting algorithm with a single execution, we must first test our concept on a real-world microcontroller. Also, a thorough assessment of the accuracy and false positives is required. Therefore, we choose three different programs, namely tinyaes128 [43], tinysha256 [45], lz4 [46], and assess our algorithm’s efficiency on them. For all three target programs, we set the threshold $\mathrm{\theta}$ to 0.3, the sampling frequency was 3.125 GS/s, and the target processor runs on an 8 MHz clock. We ran these programs on an STM32F3 board with a soldered shunt resistor. Table III illustrates the result of the experiment. Accuracy is the number of detected severe PC changes compared to the existing ones, and false positive is the rate of falsely detected changes compared to the existing ones. As the results indicate, the proposed method can identify changes in the PC and represent it as a unique fingerprint for the code execution.
Our proposed signal-processing algorithm performs well in fingerprinting the control flow of a black-box target. The most crucial point about our algorithm is that it can work with a single power trace. However, the downside of this method is the high resolution (3.125 GS/s) and high amount of computations. Also, as mentioned in Section 4.5, peripheral access can cause massive voltage peaks that may affect the accuracy of our method. Later, we evaluate our method in such cases and try to improve it.